# Few-Shot Recognition via Stage-wise Retrieval-Augmented Finetuning

### v2

@@ Abstract
Few-shot recognition (FSR) aims to train a classification model with only a few labeled examples of each concept concerned by a downstream task, where data annotation cost can be prohibitively high. We develop methods to solve FSR by leveraging a pretrained Vision-Language Model (VLM). We particularly explore retrieval-augmented learning (RAL), which retrieves data from the VLM’s pretraining set to learn better models for serving downstream tasks. RAL has been widely studied in zero-shot recognition but remains under-explored in FSR. Although applying RAL to FSR may seem straightforward, we observe interesting and novel challenges and opportunities. First, somewhat surprisingly, finetuning a VLM on a large amount of retrieved data underperforms state-of-the-art zero-shot methods. This is due to the imbalanced distribution of retrieved data and its domain gaps with the few-shot examples in the downstream task. Second, more surprisingly, we find that simply finetuning a VLM solely on few-shot examples significantly outperforms previous FSR methods, and finetuning on the mix of retrieved and few-shot data yields even better results. Third, to mitigate the imbalanced distribution and domain gap issues, we propose Stage-Wise retrieval-Augmented fineTuning (SWAT), which involves end-to-end finetuning on mixed data in the first stage and retraining the classifier on the few-shot data in the second stage. Extensive experiments on nine popular benchmarks demonstrate that SWAT significantly outperforms previous methods by $>$6% accuracy.
@@ 1 Introduction
Few-shot recognition (FSR) aims to train a model with only a few examples per concept for downstream tasks, where data annotation can be costly. One motivational application is to train machines for automated data annotation by learning from a few examples of each concept provided by data annotation guidelines [37]. Motivated by this data annotation application, we focus on solving FSR, particularly by leveraging a foundational Vision-Language Model (VLM) and its pretraining data.
@@
Motivation. Different motivations drive different FSR methods. Many existing FSR methods are factually motivated to study PEFT [52, 58, 36, 66] or learned models’ robustness and generalization [80, 81, 27]. This motivation has led to a “flawed” setup, as noted in [35, 57], where an unrealistically large validation set is used for hyperparameter tuning. In contrast, we solve FSR motivated by data annotation [37], avoiding large validation sets and prioritizing accuracy over learning a small number of parameters. This data annotation-driven perspective expands FSR research beyond the prevailing PEFT [15, 76, 35, 57, 60]. Interestingly, we find that finetuning a pretrained visual encoder on few-shot annotated examples already surpasses prior FSR methods without overfitting issues (Fig. 1)!
@@
Furthermore, most FSR methods focus on adapting a pretrained Vision-Language Model (VLM) [35, 57, 60]. Differently, we also embrace VLM’s pretraining data, exploiting retrieval-augmented learning (RAL) [33, 36, 65, 22, 50, 42]. This strategy shines in zero-shot recognition, where state-of-the-art zero-shot methods adopt a VLM to retrieve relevant pretraining data for the concepts of interest and learn a classification model over such retrieved data [42]. Our work, for the first time, extend RAL to FSR. Despite its simplicity, we reveal interesting and novel challenges and opportunities. We elaborate on them below.
@@
Challenges and Insights. Unlike prior FSR methods that learn a small number of parameters, we first explore full finetuning of the entire pretrained model (i.e., VLM’s visual encoder) on the few-shot examples. Somewhat surprisingly, this simple method significantly outperforms previous FSR methods (Fig. 1) with quite affordable computation costs (Table 5), owning to the small scale of few-shot data. In another line of research, state-of-the-art methods of zero-shot recognition (ZSR) adopt a strategy called retrieval-augmented learning (RAL), which retrieves relevant examples from VLM’s pretraining set and learns a classifier over them [42]. We test RAL for FSR by only using the retrieved data: both classifier learning (i.e., REAL-Linear [42]) and full finetuning barely surpass the non-RAL ZSR method REAL-Prompt [42] and underperform other FSR methods [57, 35], even trained with far more retrieved data. We identify two culprits (Fig. 3): (1) imbalanced distribution of retrieved data, and (2) domain gaps between the retrieved data and few-shot examples. To address the two issues, we propose our method, Stage-Wise retrieval-Augmented fineTuning (SWAT, cf. Fig. 2). SWAT first finetunes the VLM’s visual encoder on mixed retrieved and few-shot data, then re-trains the classifier solely on few-shot examples. This stage-wise training of our SWAT aligns with the decoupling learning strategy [26], which addresses the imbalanced distribution issue by first learning feature representations using all imbalanced data and then learning the classifier with balancing techniques. SWAT also mitigates the domain gap in a transfer learning manner, i.e., pretraining on larger source-domain data followed by finetuning on the target-domain data [34, 21, 19].
@@
Figure 2: Overview of our Stage-Wise retrieval-Augmented fine-Tuning (SWAT) for few-shot recognition (FSR). Consider the scenario where one wants to train a model on a few examples per concept concerned in data annotation guidelines. SWAT exploits a pretrained Vision-Language Model (VLM) and retrieves its pretraining data relevant to the concepts of interest. We observe that the retrieved data follows an imbalanced distribution and has domain gaps from the few-shot examples (Fig. 3). SWAT addresses the two issues jointly by first finetuning the VLM’s visual encoder on both retrieved and few-shot annotated data, then re-training the classifier only using the few-shot examples. Over nine FSR benchmarks, our SWAT achieves state-of-the-art performance, significantly outperforming previous methods by $>$6% accuracy (Fig. 1).
@@ 2 Related Work
Vision-Language Models (VLMs) and Retrieval-Augmented Learning (RAL). VLMs, such as CLIP [47] and ALIGN [23], are pre-trained on web-scale data sampled from the open world. They learn to map image and text data into a common embedding space via contrastive learning on millions of image-text pairs. VLMs have demonstrated impressive zero-shot transfer capability across various downstream tasks. To further improve VLMs’ zero-shot accuracy, recent works [42, 36, 65, 22, 33] explore RAL, which retrieve external data (e.g., from the VLM’s pretraining dataset) relevant to the downstream task and then use such data to finetune the VLM or learn a classifier. Inspired by RAL’s success in achieving state-of-the-art zero-shot performance, for the first time, we extend RAL to few-shot recognition. We identify not only opportunities but also interesting challenges. We present a novel FSR method that addresses these challenges, achieving state-of-the-art few-shot recognition performance on standard benchmark datasets.
@@
Moreover, to tackle data scarcity and enhance model performance, data augmentation is widely adopted during model training [49, 56]. Some augmentation techniques are performed on individual data examples, e.g., random cropping and color jittering [9, 28, 64], while others mix multiple examples [6], e.g., MixUp [75] and CutMix [74]. Among various data augmentation techniques, CutMix proves to stabilize training, reduce overfitting, and improve model performance [74]. Later works extend CutMix by combining it with MixUp [43], smoothing patch boundaries [43], leveraging saliency masks [62, 29], and considering long-tailed distribution of training data [44]. Our approach augments data by mixing retrieved images with few-shot annotated ones, yielding substantial improvements in FSR.
Figure 3: Retrieved data shows domain gaps with downstream few-shot data and follows an imbalanced distribution. Left: we compare retrieved and few-shot annotated images for random categories from five benchmark datasets. The two sets of images exhibit clear domain gaps regarding image styles, background content, and even semantics, e.g., the animal with banded stripes in the DTD dataset. Right: retrieved data follows imbalanced distributions w.r.t concepts defined in different downstream tasks, as the VLM’s pretraining set does not contain sufficient examples for certain classes. Due to these two issues, leveraging the retrieved data to improve FSR presents significant challenges. Refer to Table 1 for quantitative justification of domain gaps, and Appendix Fig. 7 and 8 for additional examples of more datasets.
3 Our Proposed Methods
@@
Problem Setup. Our FSR setup aims for high recognition accuracy on the concepts concerned by a downstream task. Our setup follows recent studies [77, 35, 57] that leverage Vision-Language Models (VLMs) for FSR. This setup also permits the use of VLM pretraining data [54, 53], available as “free lunch”, which has been exploited in zero-shot recognition [42, 36, 65, 22, 33] and is thus applicable to FSR. To support real-world applications like data annotation, this setup prioritizes recognition accuracy over limiting the number of parameter to learn, contrasting with many FSR works that treat FSR as a proxy for studying model robustness or generalization [27, 80, 81], and PEFT methods [52, 58, 36, 66].
@@
Remarks. Previous FSR methods often evaluate the generalization of models in a base-to-novel setup, which splits an existing dataset into separate base and novel class sets. In this setup, models train on few-shot data from the “base classes” and are then evaluated on the “novel classes”. However, this base-to-novel setup does not align with the our motivation – the practical application of data annotation, where the annotation guidelines clearly define the target concepts of interest, without introducing extra novel classes. More importantly, as noted by [37], this base-to-novel setup appears to be artificial as VLMs have already seen examples of these “novel” classes in their Internet-scale pretraining dataset. Therefore, in this work, we follow [37] to evaluate FSR methods on the target classes concerned by a downstream task, without a base-novel split.
Table 1: We demonstrate the domain gaps between two data sources: retrieved data and few-shot annotated data from a downstream task. Following [61], we train a binary classification model on the two data sources to identify the data source of each image from a held-out testing set. We report the binary classification accuracy on the testing sets, which are created specifically for this study based on each of downstream dataset. Results validate the presence of domain gaps with $>$90% mean accuracy over different downstream-task datasets. Refer to Fig. 3 for a visual demonstration.
@@ 3.1 Finetuning on Few-Shot Data
Recall that our FSR setup prioritizes high recognition accuracy for downstream tasks, different from prior works that focus on PEFT [35, 57]. This motivated us to first test finetuning the VLM’s visual encoder on the few-shot examples alone. To the best of our knowledge, this method has not been explored in the literature. Surprisingly, this embarrassingly simple method significantly outperforms previous FSR approaches by $>$3% (Fig. 1 and Table 2)! Although one may suspect overfitting when finetuning the whole visual encoder on the few-shot examples, our empirical results show that overfitting is not an issue (Table 2 and Fig. 4). This is likely because of the robust hyperparameters derived from other research lines [42, 35, 67, 32], including classifier initialization using text embedding (Table 12) and a smaller learning rate for updating the visual encoder (Appendix Section B). Moreover, owning to the small scale of few-shot data, learning more parameters (i.e., finetuning the entire pretrained visual encoder) is still quite efficient (Table 5).
@@ 3.2 Finetuning on Retrieved Data
Retrieval Augmentation. RAL has proven effective in zero-shot recognition [36, 65, 42]. There exist various strategies for retrieving VLM’s pretraining data relevant to concepts of interest. Some methods match feature similarities between textual embeddings of target concepts and features of pretraining images or captions [36, 65], while others use string matching to find relevant text and retrieve corresponding images [42]. The latter approach is significantly more efficient, as it avoids the high storage costs of downloading all images and the computational expense of embedding calculations for each image. Importantly, the latter helps achieve better zero-shot recognition and is more efficient than the former [42]. Therefore, we adopt the string matching approach [42] to retrieve examples in this paper.
@@ 3.3 Stage-Wise Retrieval-Augmented Finetuning
stage-1: finetuning stage-2: classifier retraining
@@ 4.1 Experimental Setup
Datasets and Metrics. Motivated from the data annotation perspective, we study FSR through challenging real-world tasks where the data annotation requires domain expert knowledge, such as recognizing bird species, aircraft models, satellite images, etc. Specifically, we follow [13, 50] to use fine-grained recognition datasets where the VLM CLIP [47] struggles to recognize the nuanced attributes [50]. We select nine datasets, namely Semi-Aves [59], Flowers102 [40], FGVC-Aircraft [38], EuroSAT [38], DTD [11], OxfordPets [45], Food101 [4], StanfordCars [31], and ImageNet [12] (see detailed descriptions in Appendix Table 6). We evaluate SWAT using the OpenCLIP ViT-B/32 model [20] and retrieve images from its publicly available LAION-400M pretraining set [53, 54]. We report the accuracy averaged over nine datasets. The appendix includes results using other VLM architectures, on which our conclusions still hold.
Table 4: Ablation study on important components in our SWAT. Comparing to the state-of-the-art adapter-based few-shot method CLAP [57], our few-shot finetuning (FSFT) obtains $>$2% improvement! Naively adding retrieved data for finetuning brings small improvements $\sim$1due to the domain gaps and imbalanced distribution of the retrieved data. Applying CutMix provides an additional $\sim$2improvement, and retraining the classifier in the second stage brings a further $\sim$0.5increase. Note the improvements of classifier retraining is more significant for the rare classes ($>$2%), especially when the stage-1 finetuning data is imbalanced retrieved data ($>$11%), as shown in Table 3. Superscripts indicate the improvements of each component (relative to the corresponding row above). We report mean accuracies averaged over nine benchmark datasets. Detailed results are listed in Appendix Table 14.
@@
Compared Methods. We compare SWAT with state-of-the-art zero-shot and few-shot recognition methods. For zero-shot methods, we study OpenCLIP [20] and the recent methods: REAL-Prompt [42] and REAL-Linear [42]. For few-shot methods, we compare against prompt-learning-based methods CoOp [80] and PLOT [8], and adapter-based methods CLIP-Adapter [15], TIP-Adapter [76], TaskRes [73], Cross-modal Linear Probing [35] and CLAP [57]. Note that CLAP is the state-of-the-art FSR method that strictly adopts the realistic validation-free setup [57].
Implementations. For each dataset, we retrieve 500 images per class following [42]. We strictly follow the validation-free protocol as done in [57] and set the same hyperparameters in all datasets (see details in Appendix Section B). We conduct experiments using 4, 8, and 16 shots of data, randomly sampled from downstream training sets with three different seeds. We report their average accuracy. Note that in the motivational application, data annotation guidelines provide multiple visual examples. Therefore, we do not conduct experiments with fewer shots (e.g., 1 or 2-shot) in our experiments. We run all experiments on a Quadro RTX 6000 (24GB) GPU with 50GB storage for hosting retrieved data for all datasets. SWAT takes 2.5 hours for larger datasets like Semi-Aves (200 classes) and less than 4 minutes for smaller ones like EuroSAT (10 classes). Table 5 compares its compute cost with previous methods. We provide our code for open source in the supplementary material.

### v3

@@ Abstract
Few-shot recognition (FSR) aims to train a classification model with only a few labeled examples of each concept concerned by a downstream task, where data annotation cost can be prohibitively high. We develop methods to solve FSR by leveraging a pretrained Vision-Language Model (VLM). We particularly explore retrieval-augmented learning (RAL), which retrieves open data, e.g., the VLM’s pretraining dataset, to learn models for better serving downstream tasks. RAL has been studied in zero-shot recognition but remains under-explored in FSR. Although applying RAL to FSR may seem straightforward, we observe interesting and novel challenges and opportunities. First, somewhat surprisingly, finetuning a VLM on a large amount of retrieved data underperforms state-of-the-art zero-shot methods. This is due to the imbalanced distribution of retrieved data and its domain gaps with the few-shot examples in the downstream task. Second, more surprisingly, we find that simply finetuning a VLM solely on few-shot examples significantly outperforms previous FSR methods, and finetuning on the mix of retrieved and few-shot data yields even better results. Third, to mitigate the imbalanced distribution and domain gap issues, we propose Stage-Wise retrieval-Augmented fineTuning (SWAT), which involves end-to-end finetuning on mixed data in the first stage and retraining the classifier on the few-shot data in the second stage. Extensive experiments on nine popular benchmarks demonstrate that SWAT significantly outperforms previous methods by $>$6% accuracy.
@@ 1 Introduction
Few-shot recognition (FSR) aims to train a model with only a few examples per concept provided by a downstream task, where data annotation can be costly. One motivational application is to train machines for automated data annotation by learning from a few examples of each concept provided by data annotation guidelines. Motivated by this data annotation application, we focus on solving FSR, particularly by leveraging a foundational Vision-Language Model (VLM) and its web-scale pretraining data.
Figure 2: Overview of our Stage-Wise retrieval-Augmented fine-Tuning (SWAT) for few-shot recognition (FSR). Consider the scenario where one wants to train a model on a few examples per concept concerned in data annotation guidelines. SWAT exploits a pretrained Vision-Language Model (VLM) and retrieves open data, e.g., the VLM’s pretraining data relevant to the concepts of interest. We observe that the retrieved data follows an imbalanced distribution and has domain gaps from the few-shot examples (Fig. 3). SWAT addresses the two issues jointly by first end-to-end finetuning the VLM’s visual encoder on mixed retrieved and few-shot annotated data, then re-training the classifier only using the few-shot examples. Over nine FSR benchmarks, our SWAT achieves state-of-the-art performance, significantly outperforming previous methods by $>$6% accuracy (Fig. 1).
@@
Motivation. Different motivations drive different FSR methods. Many existing FSR methods are factually motivated to study PEFT [54, 60, 37, 68] or learned models’ robustness and generalization [83, 84, 28]. This motivation has led to a flawed setup, as noted in [36, 59], where an unrealistically large validation set is used for hyperparameter tuning. In contrast, we solve FSR motivated by data annotation [38], avoiding large validation sets and prioritizing accuracy over learning a small number of parameters. This data annotation perspective expands FSR research beyond the prevailing PEFT [15, 78, 36, 59, 62]. Interestingly, we find that finetuning a pretrained visual encoder on few-shot annotated examples already surpasses prior FSR methods without overfitting issues (Fig. 1)!
@@
Furthermore, most FSR methods focus on adapting an open-source pretrained Vision-Language Model (VLM) [36, 59, 62]. Differently, we also embrace VLM’s pretraining data as open data, exploiting retrieval-augmented learning (RAL) [34, 37, 67, 23, 52, 43] to address the challenge of limited labeled data. RAL shines in zero-shot recognition, where state-of-the-art zero-shot methods adopt a VLM to retrieve relevant pretraining data for the concepts of interest and learn a classification model over such retrieved data [43]. Our work, for the first time, extends RAL to FSR. Despite its simplicity, we reveal interesting and novel challenges and opportunities. We elaborate on them below.
@@
Challenges and Insights. Unlike prior FSR methods that learn a small number of parameters, we first explore full finetuning of the entire pretrained model (i.e., VLM’s visual encoder) on the few-shot examples. Somewhat surprisingly, this simple method significantly outperforms previous FSR methods (Fig. 1) with quite affordable computation costs (Table 5), owning to the small scale of few-shot data. In another line of research, state-of-the-art methods of zero-shot recognition (ZSR) adopt a strategy called retrieval-augmented learning (RAL), which retrieves relevant examples from VLM’s pretraining set and learns a classifier over them [43]. We test RAL for FSR by only using the retrieved data: both classifier learning (i.e., REAL-Linear [43]) and full finetuning barely surpass the non-RAL ZSR method REAL-Prompt [43] and underperform other FSR methods [59, 36], even trained with far more retrieved data. We identify two culprits (Fig. 3): (1) imbalanced distribution of retrieved data, and (2) domain gaps between the retrieved data and few-shot examples. To address the two issues, we propose a simple method, Stage-Wise retrieval-Augmented fineTuning (SWAT, cf. Fig. 2). SWAT first finetunes the VLM’s visual encoder on mixed retrieved and few-shot data, then re-trains the classifier solely on few-shot examples. This stage-wise training of our SWAT aligns with the decoupling learning strategy [27], which addresses the imbalanced distribution issue by first learning feature representations using all imbalanced data and then learning the classifier with balancing techniques. SWAT also mitigates the domain gap in a transfer learning manner, i.e., pretraining on larger source-domain data followed by finetuning on the target-domain data [35, 22, 20].
@@ 2 Related Work
Vision-Language Models (VLMs) and Retrieval-Augmented Learning (RAL). VLMs, such as CLIP [48] and ALIGN [24], are pre-trained on web-scale data sampled from the open world. They learn to map image and text data into a common embedding space via contrastive learning on millions of image-text pairs. VLMs have demonstrated impressive zero-shot transfer capability across various downstream tasks. To further improve VLMs’ zero-shot accuracy, recent works [43, 37, 67, 23, 34] explore RAL, which retrieves open data (e.g., from the VLM’s pretraining dataset) relevant to the downstream task and then uses such data to finetune the VLM or learn a classifier. Inspired by RAL’s success in achieving state-of-the-art zero-shot performance, for the first time, we extend RAL to few-shot recognition. We identify not only opportunities but also interesting challenges. We present a novel FSR method that addresses these challenges, achieving state-of-the-art few-shot recognition performance on standard benchmark datasets.
@@
Moreover, to tackle data scarcity and enhance model performance, data augmentation is widely adopted during model training [58, 81, 51]. Some augmentation techniques are performed on individual data examples, e.g., random cropping and color jittering [9, 29, 66], while others mix multiple examples [6], e.g., MixUp [77] and CutMix [76]. Among various data augmentation techniques, CutMix proves to stabilize training, reduce overfitting, and improve model performance [76]. Later works extend CutMix by combining it with MixUp [44], smoothing patch boundaries [44], leveraging saliency masks [64, 30], and considering long-tailed distribution of training data [45]. Our approach augments data by mixing retrieved images with few-shot annotated ones, yielding substantial improvements in FSR.
Figure 3: Retrieved data shows domain gaps with downstream few-shot data and follows an imbalanced distribution. Left: we compare retrieved and few-shot annotated images for random categories from five benchmark datasets. The two sets of images exhibit clear domain gaps regarding image styles, background content, and even semantics, e.g., the animal with banded stripes in the DTD dataset. Right: retrieved data follows imbalanced distributions w.r.t concepts defined in different downstream tasks, as the VLM’s pretraining set does not contain sufficient examples for certain classes. Due to these two issues, leveraging the retrieved data to improve FSR presents significant challenges. Refer to Table 1 for quantitative justification of domain gaps, and Appendix Fig. 10 and 11 for additional examples of more datasets.
3 Problem Formulation and Methods
@@
Problem Setup. Our FSR setup aims for high recognition accuracy on the concepts concerned by a downstream task. Our setup follows recent studies [79, 36, 59] that leverage Vision-Language Models (VLMs) for FSR. This setup also permits the use of VLM pretraining data [56, 55], available as “free lunch”, which has been exploited in zero-shot recognition [43, 37, 67, 23, 34] and is thus applicable to FSR. To support real-world applications like data annotation, this setup prioritizes recognition accuracy over limiting the number of parameters to learn, contrasting with many FSR works that treat FSR as a proxy for studying model robustness or generalization [28, 83, 84], and PEFT methods [54, 60, 37, 68].
@@
Remarks. Previous FSR methods often evaluate the generalization of models in a base-to-novel setup, which artificially splits an existing dataset into separate base and novel class sets. In this setup, models train on few-shot data from the “base classes” and are then evaluated on the “novel classes”. However, this base-to-novel setup does not align with our motivation – the practical application of data annotation, where the annotation guidelines clearly define the target concepts of interest, without introducing extra novel classes. More importantly, as noted by [38], this base-to-novel setup appears to be artificial as VLMs have already seen examples of these “novel” classes in their Internet-scale pretraining dataset. Therefore, in this work, we follow [38] to evaluate FSR methods on the target classes concerned by a downstream task, without a base-novel split.
Table 1: We demonstrate the domain gaps between two data sources: retrieved data and few-shot annotated data from a downstream task. Following [63], we train a binary classification model on the two data sources to identify the data source of each image from a held-out testing set. We report the binary classification accuracy on the testing sets, which are created specifically for this study based on each downstream dataset. Results validate the presence of domain gaps with $>$90% mean accuracy over different downstream-task datasets. Refer to Fig. 3 for a visual demonstration.
@@ 3.1 Finetuning on Few-Shot Data
Recall that our FSR setup prioritizes high recognition accuracy for downstream tasks, different from prior works that focus on PEFT [36, 59]. This motivated us to first test few-shot finetuning (FSFT), which finetunes the VLM’s visual encoder only on the few-shot examples. To the best of our knowledge, this method has not been explored in the literature. Surprisingly, this embarrassingly simple method significantly outperforms previous FSR approaches by $>$3% (Fig. 1 and Table 2)! Although one may suspect overfitting when finetuning the whole visual encoder on the few-shot examples, our empirical results show that overfitting is not an issue (Table 2 and Fig. 4). This is likely because of the robust hyperparameters derived from other research lines [43, 36, 69, 33], including classifier initialization using text embedding (Table 14) and a smaller learning rate for updating the visual encoder (Appendix Section B). Moreover, owning to the small scale of few-shot data, learning more parameters (i.e., finetuning the entire pretrained visual encoder) is still quite efficient (Table 5).
@@ 3.2 Finetuning on Retrieved Data
Retrieval Augmentation. RAL has proven effective in zero-shot recognition [37, 67, 43]. There are various strategies for retrieving open data, especially VLM’s pretraining data, relevant to concepts of interest. Some methods match feature similarities between textual embeddings of target concepts and features of pretraining images or captions [37, 67], while others use string matching to find relevant text and retrieve corresponding images [43]. The latter approach is significantly more efficient, as it avoids the high storage costs of downloading all images and the computational expense of embedding calculations for each image. Importantly, the latter helps achieve better zero-shot recognition and is more efficient than the former [43]. Therefore, we adopt the string matching approach [43] to retrieve examples in this paper.
@@ 3.3 Stage-Wise Retrieval-Augmented Finetuning
stage-1: finetuning stage-2: classifier retraining on few-shot data
@@ 4.1 Experimental Setup
Datasets and Metrics. Motivated from the data annotation perspective, we study FSR through challenging real-world tasks where the data annotation requires domain expert knowledge, such as recognizing bird species, aircraft models, satellite images, etc. Specifically, we follow [13, 52] to use fine-grained recognition datasets where the VLM CLIP [48] struggles to recognize the nuanced attributes [52]. We select nine datasets, namely Semi-Aves [61], Flowers102 [41], FGVC-Aircraft [39], EuroSAT [39], DTD [11], OxfordPets [46], Food101 [4], StanfordCars [32], and ImageNet [12] (see detailed descriptions in Appendix Table 6). We evaluate SWAT using the OpenCLIP ViT-B/32 model [21] (unless otherwise specified) and retrieve images from its publicly available LAION-400M pretraining set [55, 56]. We report the accuracy averaged over nine datasets. The appendix includes results using other VLM architectures, on which our conclusions still hold.
Table 4: Ablation study on important components in our SWAT. Compared to the state-of-the-art adapter-based few-shot method CLAP [59], our few-shot finetuning (FSFT) obtains $>$2% improvement! Naively adding retrieved data for finetuning brings small improvements $\sim$1due to the domain gaps and imbalanced distribution of the retrieved data. Applying CutMix provides an additional $\sim$2improvement, and retraining the classifier in the second stage brings a further $\sim$0.5increase. Note the improvements of classifier retraining are more significant for the rare classes ($>$2%), especially when the stage-1 finetuning data is imbalanced retrieved data ($>$11%), as shown in Table 3. Superscripts indicate the improvements of each component (relative to the corresponding row above). We report mean accuracies averaged over nine benchmark datasets. Detailed results are listed in Appendix Table 16.
@@
Compared Methods. We compare SWAT with state-of-the-art zero-shot and few-shot recognition methods. For zero-shot methods, we study OpenCLIP [21] and the recent methods: REAL-Prompt [43] and REAL-Linear [43]. For few-shot methods, we compare against prompt-learning-based methods CoOp [83] and PLOT [8], and adapter-based methods CLIP-Adapter [15], TIP-Adapter [78], TaskRes [75], Cross-modal Linear Probing [36] and CLAP [59]. Note that CLAP is the state-of-the-art FSR method that strictly adopts the realistic validation-free setup [59]. We present the results of different finetuning methods in Table 13 of the Appendix.
Implementations. For each dataset, we retrieve 500 images per class following [43]. We strictly follow the validation-free protocol as done in [59] and set the same hyperparameters in all datasets (see details in Appendix Section B). We conduct experiments using 4, 8, and 16 shots of data, randomly sampled from downstream training sets with three different seeds. We report their average accuracy. Note that in the motivational application, data annotation guidelines provide multiple visual examples. Therefore, we do not conduct experiments with fewer shots (e.g., 1 or 2-shot) in our experiments. We run all experiments on a Quadro RTX 6000 (24GB) GPU with 50GB storage for hosting retrieved data for all datasets. SWAT takes 2.5 hours for larger datasets like Semi-Aves (200 classes) and less than 4 minutes for smaller ones like EuroSAT (10 classes). Table 5 compares its compute cost with previous methods. We provide our code instructions in Appendix Section I.